OpenStudy (anonymous):

A jugler tosses three balls alternately uoward . Eaxh ball has an initial velocity of 5m/s how high does each ball rise?

OpenStudy (anonymous):

\(v_i=5 m/s, v_f=0 m/s, a=9.8m/s^2, d=?\) \(v_f^2=v_i^2+2ad\) rearrange this kinematics equation to solve for "d". \(\Large d=\frac{v_f^2-v_i^2}{2a}\) substitute your given values and you're done :)